OpenStudy (anonymous):

Find the limit if it exists, or show that the limit does not exist. lim (x^2+y^2)/(sqrt(x^2+y^2+1)-1) (x,y)->(0,0) I've tried approaching the limit from the line y = 0 and y = x. Then I thought I could apply the squeeze theorem. No luck. The book says the answer is 2

OpenStudy (anonymous):

\[\lim_{(x,y) \rightarrow (0,0)} \frac{ x^2+y^2 }{ \sqrt{ x^2+y^2+1}-1 }\] Sorry about changing the answer so many times. Devil is in the details.

OpenStudy (kenljw):

you end up with the result 0/0 use L'Hospital Rule lim (x,y) to (0,0) 2x_2yy'/(1/2 sqrt(x^2 +y^2+1)^-1/2(2x+2yy') lim (x,y) to (0,0) 2sqrt(x^2+y^2+1) = 1 been awhile check

OpenStudy (kenljw):

2 not 1

OpenStudy (anonymous):

Sorry for my ignorance, I haven't taken calculus in 6 years. From what you wrote I should differentiate the equation with respect to y?

OpenStudy (kenljw):

yes using implicit differentiation

OpenStudy (anonymous):

I'll get on that. Thank you.

OpenStudy (experimentx):

try switching coordinates ,,, use polar coordinates

OpenStudy (experimentx):

also try switching 1/x and 1/y and change limits to infinity and use this theorem http://math.stackexchange.com/questions/15240/when-can-you-switch-the-order-of-limits

OpenStudy (anonymous):

Thanks. I'm currently trying to work through KenLJW's suggestion and make sure I'm doing everything correctly. After that I'll look up polar coordinates again and try that approach.

OpenStudy (experimentx):

it's a double sequence, you can'f differentiate 'y' w.r.t. 'x', they are independent coordinates ... dy/dx is zero naturally.

OpenStudy (anonymous):

The approaches we've covered to this point are l'hopital's, squeeze, and approaching from different lines. So far I'm working with l'hopital to try and get the correct answer. This question is not assigned and it's not mentioned under the suggested problems. It's just in between suggested problems so I assumed I had the tools to complete it.

OpenStudy (experimentx):

use approaching from different lines then, put y=mx, and see if you get same limit for different values of 'm'. if limit is same then the limit exists. If limit is different then it does not exist.

OpenStudy (experimentx):

perhaps you can use L'hopital rule after that.

OpenStudy (anonymous):

At this point my understanding is that there could always exist another line such that the limit is different because we can approach the point (0, 0) from an infinite number of possibilities.

OpenStudy (anonymous):

I'll try y=mx with l'hopital's rule.

OpenStudy (experimentx):

good luck!! I gotta sleep

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

So Ken, I'm unable to get the same values as you. From your post I read the intermediate step I should get is \[\lim_{(x,y) \rightarrow (0,0)} \frac{2xyy'1/2 } { \sqrt(x^2+y^2+1)^{-\frac{1}{2}(2x+2yy')}}\], is this correct?

OpenStudy (kenljw):

The way I did it the Numerator canceled?

OpenStudy (kenljw):

N' = 2x+2yy' D' = 1/2 sqrt(x^2+y^2+1)^-1/2 (2x+2yy')

OpenStudy (anonymous):

I still don't get that. Using the product rule I get a massive equation.

OpenStudy (anonymous):

I think I'll bring this into office hours tomorrow. Thanks for your help guys.

OpenStudy (kenljw):

It's not the product rule it's the chain rule in the denominator

OpenStudy (anonymous):

Someone from class showed me how to solve it through substitution. Thanks for all your help! \[\lim_{\left(x,y\right) \rightarrow \left(0,0\right)} \frac{x^2+y^2}{\sqrt{x^2+y^2+1}-1} \] \[v=\sqrt{x^2+y^2+1}\] \[\lim_{\left(x,y\right) \rightarrow \left(0,0\right)} \frac{a^2-1}{a-1}\] \[u=a-1\] \[\lim_{\left(x,y\right) \rightarrow \left(0,0\right)} \frac{\left(u+1\right)^2-1}{u}\] \[\lim_{\left(x,y\right) \rightarrow \left(0,0\right)} \frac{u+2}{u} = \lim_{\left(x,y\right) \rightarrow \left(0,0\right)} u+2\] \[\lim_{\left(x,y\right) \rightarrow \left(0,0\right)} u+2 = \lim_{\left(x,y\right) \rightarrow \left(0,0\right)} a -1 + 2\] \[\lim_{\left(x,y\right) \rightarrow \left(0,0\right)} \sqrt{x^2+y^2+1}+1 = 2\]

OpenStudy (anonymous):

typo in the above response. v = a

OpenStudy (kenljw):

Same as my answer which was done direct using chain rule

OpenStudy (experimentx):

that's same as using using polar coordinates http://www.wolframalpha.com/input/?i=limit+r-%3E0+r%5E2%2F%28sqrt%28r%5E2+%2B+1%29+-+1%29 although that substitution is clever than polar

OpenStudy (kenljw):

Even with polar you still have to use L'Hospital and change rule if you don't graph

OpenStudy (kenljw):

When you made substitution all you had to do was factor numerator then cancel giving direct answer

OpenStudy (experimentx):

yes that it is, but I am not sure of method suggested. I've never seen using differentiation using on double limits

OpenStudy (kenljw):

It's okay if x goes to 0 then y go to 0 is the same as y goes to 0 then s goes to 0

OpenStudy (kenljw):

from either direction

OpenStudy (experimentx):

f(x,y) = xy/(x^2 + y^4) limit (x,y) = (0,0) this case it's different.

OpenStudy (kenljw):

f(x,y)=yx^3+xy^3 an apparent odd function? limit close by is not symmetric?

OpenStudy (experimentx):

(x,y)->(0,0) your limit is zero .. i guess

OpenStudy (kenljw):

I'm going on the assumption close to 0 must be symmetric near by. why don't you graph it on your Mathematica program

OpenStudy (experimentx):

the function is continuous at (x,y) = (0,0), naturally limit is zero.